Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00 per unit, and fixed expenses total $160,000 per year. Its operating results for last year were as follows:

Sales $ 2,080,000
Variable expenses 1,040,000
Contribution margin 1,040,000
Fixed expenses 160,000
Net operating income $ 880,000

Required:

Answer each question independently based on the original data:

1. What is the product's CM ratio?

2. Use the CM ratio to determine the break-even point in dollar sales.

3. If this year's sales increase by $59,000 and fixed expenses do not change, how much will net operating income increase?

4-a. What is the degree of operating leverage based on last year's sales?

4-b. Assume the president expects this year's sales to increase by 20%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

5. The sales manager is convinced that a 14% reduction in the selling price, combined with a $64,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. Do you recommend implementing the sales manager's suggestions?

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.40 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $880,000 net operating income as last year?
Answer #1

1.

Contribution margin ratio = Contribution margin/Sales

= 1,040,000/2,080,000

= 50%

2.

Break even point = Fixed expenses/Contribution margin ratio

= 160,000/50%

= $320,000

3.

Increase in sales = $59,000

Hence, this year's sales = 2,080,000 + 59,000

= $2,139,000

Net operating income = Sales x Contribution margin ratio - Fixed expenses

= 2,139,000 x 50% - 160,000

= $909,500

4-a

Degree of operating leverage = Contribution margin/Operating income

= 1,040,000/880,000

= 1.1818

4-b

Increase in sales = 20%

Degree of operating leverage = % change in operating income/% change in sales

1.1818 = % increase in operating income/20

Hence, % increase in operating income = 23.63%

Due to 20% increase in sales, operating income of this year will increase by 23.63%

Note: I have answered first 4 questions as per HOMEWORKLIB POLICY. Please post your other questions separately. Thanks.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.